### v1

# Mixed Density Diffuser: Efficient Planning with Non-uniform Temporal Resolution

Abstract
Recent studies demonstrate that diffusion planners benefit from sparse-step planning over single-step planning. Training models to skip steps in their trajectories helps capture long-term dependencies without additional or memory computational cost. However, predicting excessively sparse plans degrades performance. We hypothesize this temporal density threshold is non-uniform across a temporal horizon and that certain parts of a planned trajectory should be more densely planned. We propose Mixed Density Diffuser (MDD), a diffusion planner where the densities throughout the horizon are tunable hyperparameters. MDD achieves a new SOTA across the Maze2D, Franka Kitchen, and Antmaze D4RL task domains.
1 Introduction
Training a policy with online rollouts can be costly, dangerous, and sample-inefficient [1]. Alternatively, offline reinforcement learning (RL) involves a policy trained exclusively with pre-collected data. Extracting effective polices without environment exploration or feedback proves challenging for conventional off-policy and even specialized offline RL algorithms [2, 3]. A frequently occurring property of offline RL problems are incomplete, or undirected demonstrations [4, 5, 6]. Offline algorithms must compose sub-trajectories from training data to generate advantageous behaviors. Another challenge is high-dimensionality and long-horizons which makes accurate planning and behavior cloning difficult [1]. Finally, sparse rewards challenge many training algorithms as they hinder accurate action credit assignment.[7].
Diffusion probabilistic modeling is a powerful framework for expressing complex, multi-modal distributions [8, 9]. Leveraging this model class, diffusion policies generate high fidelity actions and use a value function for high-reward action selection [10, 11, 12]. Diffusion planners approach offline RL as a guided sequence generation task [13, 14, 15]. Planners predicting low temporal density, sparse-step trajectories are a promising method of lengthening temporal horizons while maintaining manageable state-space output dimensionality [16]. However, excessively low density planning degrades performance. Especially in tasks with urgent, short-term requirements, like locomotion [17]. This suggests a fundamental limitation of uniform density trajectories which balance temporal resolution and detail with horizon length. We theorize that some parts of a planned trajectory require less resolution than others [18]. Since uniform-density planners are unable to accommodate these shifts in redundant information, they often under-plan or model superfluous states.
Hierarchical planners circumvent this issue by training an ensemble of models with varying temporal resolutions and horizon lengths [16, 19]. In hierarchical approaches, a low-density planner creates coarse, long-horizon trajectories. Often as subgoals, this long horizon information is passed to a higher-density model which interpolates between the coarse states. Hierarchical models can achieve impressive performance with competitive inference and training times, but their ensemble structure means multiplying memory requirements and the number of trained parameters. The inter-dependence of a hierarchy hinders robustness by compounding errors between models. Also, using a hierarchal framework makes end-to-end planner training impossible, potentially inhibiting optimization and information flow.
To overcome the limitations of both uniform-density and hierarchical planners, we introduce Mixed Density Diffuser (MDD). Our primary contributions are:
  • MDD generates trajectories with non-uniform temporal densities using a single, flat diffusion model. Our implementation integrates mixed-density training with the powerful Diffusion Veteran (DV) framework [17].
  • MDD achieves state-of-the-art performance on several D4RL tasks without increasing model parameters or inference costs from the base DV framework.
  • We present strong empirical evidence supporting our theory that non-uniform temporal horizons are a key principle of efficient, effective, simple diffusion planning.
Figure 1: Generalized comparison of trajectory denoising in different planning densities. Sparse Density Diffusers in row $1)$ extend temporal planning horizons with comparatively little computational cost at the price of low temporal resolution. High Density Diffusers in row $2)$ compute many steps for shorter temporal horizons creating more continuous planned trajectories. Hierarchical Planners in row $3)$ [16, 18, 19] generate way-points with a Sparse Density Diffuser and interpolate between with a High-Density Diffuser. Our approach (bottom row) utilizes the benefits of sparse and dense planning with a simple, flat framework by treating the sparse and dense observations as a single trajectory for one planner to generate. Row $5)$ is fully denoised example trajectories that would be non-trivial to generate with Hierarchical Planners, but simple to implement with MDD.
2 Diffusion Planning
Diffusion planners [13, 14, 15, 17] approach offline RL as a guided sampling problem. Given a current state $s_{t}$ or state-action pair $(s_{t},a_{t})$, diffusion planners predict an $H$ step trajectory, $\tau$, with some jump variable $K$ timestep between each planned step. The initial position $X_{t}$ is fixed to the current true observation. A generalized trajectory output where $X_{t}$ can either be $s_{t}$ or $(s_{t},a_{t})$ is,
$\mathbf{\tau}=\begin{bmatrix}X_{t}&X_{t+K}&X_{t+2K}&\dots&X_{t+HK}\end{bmatrix}$
A guidance function models the expected value of trajectories and focuses sampling toward high-reward sequences. Early implementations predicted state-action trajectories, but recent studies [14, 17] empirically show generating state-only trajectories and using an inverse dynamics model $a_{t}=h(s_{t},s_{t+k})$ yields higher performance. Unlike auto-regressive world models traditionally used in model-based RL [20, 21], these diffusion planners predict states at the trajectory level, avoiding compounding errors from one-step rollouts. In existing diffusion planners, the jump variable $K$ is constant, preventing precise temporal resolution allocation–a limitation MDD addresses.
3 Mixed Density Diffuser Framework
While our approach can be flexibly integrated with many diffusion planner designs we leverage the DV framework, the product of a comprehensive hyperparameter sweep [17]. MDD uses a DiT denoising backbone[22], MCSS guidance, and a single flat planner predicting state-only trajectories paired and a diffusion-based inverse dynamics model.
We can abstract the output trajectory as,
$\mathbf{\tau}=\begin{bmatrix}x_{t}&x_{t+K_{1}}&x_{t+K_{1}+K_{2}}&\dots&x_{t+K_{1}+K_{2}+\dots+K_{H-1}}\end{bmatrix}$
where each $K_{i}$ is an independent, tunable hyperparameter. Adjusting these $K_{i}$ jump sizes allows precise control of temporal resolution. Tuning ranges of jump indices quickly narrowed our hyperparameter search and is easier to interpret.
Table 1: Normalized score comparison of offline RL methods across 150 episode seeds on the Franka Kitchen [23], Antmaze, and Maze2D D4RL tasks [7]. The best scores per row are bolded.
Kitchen Antmaze Maze2D Method Mixed Partial avg. L.-diverse L.-play M.-diverse M.-play avg. L M. Umaze avg. BC 47.5 33.8 40.7 0.0 0.0 0.0 0.0 0.0 5 30.3 3.8 13.0 BCQ [4] 8.1 18.9 13.5 2.2 6.7 0.0 0.0 2.2 6.2 8.3 12.8 9.1 CQL [2] 51.0 49.8 50.4 61.2 53.7 15.8 14.9 36.4 12.5 5.0 5.7 7.7 IQL [3] 51.0 46.3 48.7 47.5 39.6 70.0 71.2 57.1 58.6 34.9 47.4 47.0 SfBC [12] 45.4 47.9 46.7 45.5 59.3 82.0 81.3 67.0 74.4 73.8 73.9 74.0 DQL [10] 62.6 60.5 61.6 56.6 46.4 78.6 76.6 64.6 - - - - DQL* [24] 55.1 65.5 60.3 70.6 81.3 82.6 87.3 80.5 186.8 152.0 140.6 159.8 IDQL [11] 66.5 66.7 66.6 67.9 63.5 84.8 84.5 75.2 90.1 89.5 57.9 79.2 IDQL* [24] 66.5 66.7 66.6 40.0 48.7 83.3 67.3 69.8 - - - - Diffuser [13] 52.5 55.7 54.1 27.3 17.3 2.0 6.7 13.3 123 121.5 113.9 119.5 AdaptDfsr [15] 51.8 55.5 53.7 8.7 5.3 6.0 12.0 8.0 167.9 129.9 135.1 144.3 DD [14] 75.0 56.5 65.8 0.0 0.0 4.0 8.0 3.0 - - - - HD [16] 71.7 73.3 72.5 83.6 - 88.7 - - 128.4 135.6 155.8 139.9 DfsrLite [18] 73.6 74.4 74.0 80.4 72.4 89.2 88.0 82.5 - - - - DV [17] 73.6 94.0 83.8 80.0 76.4 87.4 89.0 83.2 203.6 150.7 136.6 163.6 MDD (Ours) 75.0 99.7 87.4 80.7 77.3 87.3 88.7 83.5 206.1 154.5 138.3 166.3
4 Experimental Results
We evaluate MDD on three offline RL task domains from the widely used D4RL benchmark [7]: Maze2D, AntMaze, and Franka Kitchen [23]. In Maze2D, the agent must navigate a 2D maze using a simple ball embodiment, directly testing MDD’s long-horizon planning ability. Antmaze extends Maze2D to a higher-dimensional setting by replacing the ball with an 8-DoF quadruped. The agent receives sparse rewards, only when reaching the target location. In Franka Kitchen, the agent is trained using suboptimal or incomplete demonstrations to control a 9-DoF Franka arm for a mutli-part manipulation task. Our main results are displayed in Table  1. Despite having no significant architectural or inference procedure differences from the DV framework, MDD on average outperforms DV in each task domain.
5 Discussion
Our framework adjustments only relate to the temporal density of the output trajectory states. Similarly to the Diffusion Veteran baselines, we use the same temporal density configuration for each task in a task domain demonstrating some hyperparameter robustness. Specifically, we use a denser-to-sparser approach. However, specific density configurations perform much higher on specific subtasks. A sparse-to-dense approach actually surpassed every model we compared in the Antmaze-Diverse task–scoring $84.0$. We theorize that adding higher temporal resolution to a latter part of the output trajectory gave the model more computational resources and training feedback for some long-term dependency, but leave further investigation for future work. Although MDD is outperformed slightly by $1$ to $2$ models on $4$ subtasks, the base DV framework was outperformed by an even larger margin on half of MDD’s underperforming subtasks. This suggests a single mixed density does not always generalize well to every subtask and that MDD inherits some of DV’s shortcomings. Particular, MDD and DV achieved very high performance on the Large and Medium Maze2D tasks, but were surpassed by two baselines on the Umaze task. The Umaze environment is smaller and simpler than the Large and Medium environments suggesting that approaches like HD and DQL have superior short-term planning but lack the long-horizon foresight to achieve a higher mean score on the Maze2D tasks. In both the Kitchen and Maze2D tasks, we compare two variations of the Diffusion Veteran framework to test whether MDD’s performance increase came from simply increasing the sparsity and temporal horizon of generated DV trajectories. We adopt the notation DV-J where J is the adjusted jump size. Surprisingly, we find that increasing DV’s jump size actually significantly increases it’s Kitchen performance despite DV having already been tuned in a hyperparameter sweep of over $6,000$ models [17]. MDD still outperforms these DV-J variants on average.
Table 2: DV temporal density variations of on the Franka Kitchen and Maze2D tasks. Results are averaged over 1000 episode seeds
Method Kitchen Mixed Kitchen Partial avg. Method Maze L. Maze M. Umaze avg. DV-5 74.7 99.4 87.05 DV-17 125.9 150.8 195.3 157.33 DV-6 74.65 98.5 86.6 DV-25 130.4 153.5 199.5 161.13 MDD 75.0 99.7 87.4 MDD 206.1 151.5 138.2 166.3